OpenStudy (anonymous):

Help? The Sum Of Two Numbers Is 24. When The Second Number Is Subtracted By The First Number, The Difference Is 16. Find The Two Numbers.

OpenStudy (anonymous):

Ok so x + y = 26 and x-5=16

OpenStudy (anonymous):

x= 21 because if we subtract 5 from it its 16 so 21+y=24 whats y?

OpenStudy (anonymous):

oh, okay thanks. i got it.

OpenStudy (campbell_st):

the solution posted above is incorrect... let x be the 1st number and y be the 2nd number using what you know x + y = 24 x - y = 16 ---------- add the 2 equations this will eliminate y 2x = 40 solve for x to find y, subsitute your x answer into either equation and solve for y